# Emergent interactions lead to collective frustration in robotic matter

### v1

@@ Results
A large number of environments tend to follow similar trends, indicating that the DNNs in particles do not distinguish between these environments in terms of repulsiveness and that DNNs initially do not interpret the full complexity of the environment. At characteristic times that partially overlap with the boundaries of the temporal regimes defined on the macroscopic scale we observe branching points after which environments become distinguishable by the DNNs to increasing degrees. Therefore, over time, DNNs of particles learn to distinguish between an increasing number of environments at characteristic time points. In between these time points, the form of interactions tends to be stable over increasingly long time periods.
@@
Figure 7: To understand how interactions involve over time, we computed the time evolution of the repulsiveness, averaged over all particles. Each line shows the repulsiveness in a different environment. (a) Repulsiveness colored by the distance to the particle added to the base environments. (b) Repulsiveness for the case that a particle is added at a distance $\Delta x=1$ to the base environments. Environments are coloured by the distance to the second nearest neighbour. (c) Repulsiveness for the case that a particle is added at distance $\Delta x=1$ to the base environments. Environments are coloured by the distance to the second nearest neighbour.
@@
With this, we are now in a position to come back to the question of what is the origin of the abrupt change in the behaviour of particle as a function of the particle density and the associated decrease of rewards. Our analyses have shown both: macroscopic order in space and time, which indicates collective behaviour of particles; and highly complex, density-dependent interactions that depend on evolving particle “species”. To test the hypothesis that the abrupt change in the behaviour of particles is a consequence of self-organization, we now develop a minimal, phenomenological model that only comprises a minimal set of interactions without explicit density dependence.
To this end, we use the framework of active matter theory, and Langevin equations. We consider the position $x_{i}(t)$, and velocity $v_{i}(t)$ of a particle $i$ at a given instance of time, $t$. The time evolution of the position and velocity of this particle is then described by a set of Langevin equations of the form
@@
We first identify the shape of the interactions $K_{i}$ from our simulations. To this end, we take a coarse-grained perspective and quantify the force acting on a particle $i$ for a given average velocity field $u(x)$ and density field $\rho(x)$ averaged in a small environment around the position $x_{i}$, $K_{i}(\rho,u)$. In our simulations, interactions evolve over time and are highly complex. They cannot be easily translated into simple interaction rules. For the purpose of our argument, we identify two density-, particle, and species-independent components of interactions: First, we have shown in Fig. 7(b) that interactions between particles are predominantly repulsive. This gives rise to a contribution to the interaction term of the form $-\partialderivative*{\rho(x)}{x}\evaluated{}_{x=x_{i}}$.
@@
where $\tau_{-}$ and $\tau_{+}$ are the autocorrelation times of collisions and non-collisions, respectively. The minimisation of the autocorrelation time is empirically supported by our simulations (Fig. S1). The autocorrelation time is proportional to the time a pair of particles spend in each other’s vicinity, such that particles minimise a function that is proportional to $1/|v-u|$. This then leads to a force that is given by the derivative of this term with respect to $v$, $\operatorname{sign}(v-u)/|v-u|^{2}$. Because the correlation scales with the density, the second contribution to the interaction terms takes the form $\rho\operatorname{sign}(v-u)/|v-u|^{2}$. Taken together, in our phenomenological description, this gives an effective force term for particle $i$ of the form
$K_{i}=\zeta\rho\frac{\operatorname{sign}(v-u)}{|v-u|^{2}}-\delta\partialderivative{\rho}{x}\evaluated{}_{x=x_{i}}\,.$ (5)
Figure 8: a) Representative trajectories showing the positions of particles at the specified particle density in the last 200 episodes. The colors show the sign of average velocity of the particles in the last 1e3 time points: red for positive and blue for a negative average velocity. b) Phase portrait of Eqs. (9) and (10) for density 0.1. c) phase portrait of Eqs. (9) and (10) for density 0.5. The colors show the sign of the velocity of the test particle in the Hartree approximation. d) velocity autocorrelation time obtained from simulations of Eqs. (7) and (8). The dashed line shows the density 0.3. e) Velocity autocorrelation time obtained from simulation of the full model in Fig. 1(b).
We now define the active friction term $\gamma(v)$. Because our system is out of equilibrium, the fluctuation-dissipation theorem does not hold, such that friction is not coupled to the noise amplitude. Instead, this term comprises a part that describes directed motion and a part that describes friction. A phenomenological term that describes both is the Rayleigh-Helmholtz friction [32], which has been used to describe a variety of active matter systems. It is of the form
@@
where we have dropped the index $i$. We have non-dimensionalised the Langevin equations by rescaling time and space by $t^{\prime}=\alpha t$ and $x^{\prime}=x\sqrt{\alpha\beta}$, respectively, and removed the apostrophes to simplify the notation. The dimensionless parameters $s=\zeta\sqrt{\beta^{5}/\alpha^{3}}$ and $r=\delta\sqrt{\beta^{3}/\alpha}$ quantify the strength of interactions due to the minimisation of correlations and repulsion relative to the effect of self-propulsion, respectively. $B=D\beta/\alpha^{2}$ is the dimensionless noise strength.
In this model, the relative strength of the two interaction terms implicitly depends on the particle density. For high densities, fluctuations in the particle density are weak, such that we expect the term that minimises correlations to dominate. Intuitively, this leads to a maximisation of the difference between the particle’s velocity $v$ and the average velocity $u$, which could globally lead to a split as observed in our simulations. For low densities, the term that minimises correlations is negligible and we expect particle repulsion to be dominant.
To test this numerically, we performed stochastic simulations of Eqs. (7) and (8) using a finite difference method (see Appendix C). We indeed found that particles show global alignment at low densities and a split into groups moving in opposite direction for high densities (Fig. 8). Again, the variance of particles moving in the positive direction across simulations is smaller than expected by chance ($p=2.22e^{-16}$), suggesting a collective tuning of the sizes of both groups. These observations suggest that the abrupt change in the behavior of particles in Fig. 3(a) is a phase transition resulting from collective behavior induced by interactions between particles.
To investigate this more quantitatively, we analytically derived the coevolution of the velocity of a particle, $v$, and the average velocity of all particles $\bar{u}$. To this end, followed Ref. [33] and used a moment expansion up to second order in the average velocity to obtain a set of nonlinear differential equations,
$\dot{v}=$ $(1-v^{2})v+s\frac{\rho\operatorname{sign}{(v-\bar{u})}}{(v-\bar{u})^{2}}+r(\bar{u}-v),,$ (9)
@@
Here, $T=\langle(v-\bar{u})^{2}\rangle$ is the variance of the velocities, which we estimated from our simulations of the Langevin equations. The co-evolution of the velocity of individual particles and the average velocity is summarized in phase portraits (Fig. 8). The phase portraits show attractors that correspond to asymptotic dynamical regimes of Eqs. (7) and (8). For low densities, we find two attractors with $v=\bar{u}$ that correspond to states, where the velocity of particles aligns with the average velocity in the system. For high densities, these attractors are located horizontally on a line given by $\bar{u}=0$. Therefore, the phenomenological model predicts the evolution of individual velocities to take specific positive or negative values under the constraint of a vanishing global velocity. Because the line $\bar{u}=0$ is stable, fluctuations around it are suppressed by feedback mechanisms. This corresponds to the observed state with precisely tuned equal proportions of particles moving in either direction.
Taken together, our analysis shows that the abrupt change of the behaviour as a function of the particle density is is an emergent phenomenon resulting from interactions between particles. It is therefore a phase transition as in statistical physics systems. For high densities, interactions between particles have the macroscopic effect . In this regime, particles optimise their reward locally, without being able to predict the suboptimal consequences of their behaviour on the macroscopic scale. This resembles a learning analogon to frustration phenomena in disordered systems or the fallacy of the commons in social sciences. This analysis raises the question of whether this phase transition resembles known phase transitions in (non-equilibrium) statistical physics. To investigate this, we calculated autocorrelation functions of the velocities of individual particles using the batch means method [34], both in the Langevin simulations and in the simulations of the full model comprising intelligent agents, for different particle densities. In both cases, the autocorrelation time diverges at the density at which the phase transition occurs. Diverging correlations are characteristic of critical points, suggesting critical behaviour. A further characterisation of the phase transition requires the computation of critical exponents, which is currently not feasible for intelligent agents. It is, however, of note that the phase transition occurs at a density, at which the range covered by the repulsive interactions of each particle multiplied by number of particles equals the lattice size, that is, the effective volume fraction is equal to 1. This could indicate a relation to jamming transitions.
@@ Discussion
Taken together, we studied a paradigmatic model of robotic matter. In this model, particles in a stochastic lattice gas are endowed with deep neural networks that take the spatial environment of a particle as an input and predict transition rates. Our extensive simulations reveal a rich phenomenology, including complex emergent phenomena leading to a phase transition, the evolution of multiple particle species, temporal learning regimes, and a phenomenon akin to frustration, where particles adopt suboptimal strategies. The phase transition we observe shows signs of criticality. Although it is not possible to determine critical exponents, the behaviour close to the phase transition point shows similarities to a jamming transition.
@@
Stochastic many-particle systems as the one studied in this work differ qualitatively from systems typically studied in active matter theory: because each particle has many more intrinsic degrees of freedom than the number of particles one can typically simulate on computers, they effectively lack microscopic symmetries. While in statistical physics one is typically interested in the breaking of microscopic symmetries, here the question arises if symmetries emerge as a consequence of learning, and whether these symmetries are then broken again. Because deep neural networks of sufficient size can, in principle, approximate arbitrarily complex function interactions between particles do not necessarily follow simple patterns. They may be qualitatively different between particles and environments. This raises the question of in how far physical theories are adequate in describing the time evolution of such systems. In our work, particles indeed developed complex interactions that cannot be described by simple mathematical functions. However, a simple phenomenological theory focussing on rudimentary aspects of these interactions is able to predict a key aspect of the system, namely the existence of a density-dependent phase transition.
@@
Although our model is a drastic simplification of real robotic systems, our results show similarities to much more complex robotic systems. For example, Ref. [19] studied a hide and seek game between intelligent agents, in which agents passed through different, metastable learning regimes and interactions between agents lead to a decrease in the collected reward. Apart from interacting robots, also virtual artificial intelligence systems, like the widely-used large-language systems, are now beginning to be combined in order to achieve better performance [35]. This raises the question of how to design these systems to achieve an expected outcome. In some cases, interacting learning objectives straightforwardly translate into collective behaviour [14]. Our work implies that learning can be drastically influenced in unpredictable manners by collective behaviour on the macroscopic scale.
@@ Acknowledgements
A.A. and O.B. contributed equally. A.A., O.B., and S.R. conceptualized the work. A.A. and S.R. supervised the work; O.B. wrote the simulation code and analysed O.B. performed analytical calculations. S.R. acquired funding. O.B. and S.R. wrote the initial draft of the manuscript. All authors edited and approved of the final version of the manuscript.
@@ A Implementation of the stochastic simulations
$\sum_{i=1}^{n_{j}-1}-A_{i}\log{p_{i}},$
where $n_{j}$ is the number of transitions the particle $j$ made during the episode, and $A_{i}:=R_{i}+\gamma V_{i+1}-V_{i}$, and $p_{i},R_{i},V_{i}$ are the probability with which the particle made the transition, the reward it got and the value predicted by the critic at step $i$, respectively. To train the critic, we used the loss $\sum_{i=1}^{n_{j}-1}A_{i}^{2}/(n_{j}-1)$.
Figure S1: The average autocorrelation time of collisions over 101010 replicates. The shaded area around the curves denote the standard error. The averages are calculated at logarithmically samples time points.
@@ B Numerical solution of the Langevin equations
We used the Euler scheme to integrate the Langevin equations. We used 200200200 particles with time increment $\delta t=0.01$, and the system size was determined based on $L=200/\langle\rho\rangle$ where $\langle\rho\rangle$ is the average particle density. We used periodic boundary conditions. We set $s=1.17,r=12.50,B=0.02$, and we used a local radius of 555 to calculate the local density field $\rho(x)$ and the local velocity field $u(x)$ for each particle. We run 151515 to 242424 replicates per particle density. To calculate the gradients of the local fields, we subtracted value of the field in the right of the particle at distance 555 from the value of the field in the left of the particle at distance 555, and divided this by 555.
@@ C Non-dimensionalisation of the Langevin equation
To non-dimensionalize the Langevin equations, we rescaled time and positions as
@@ D Minimisation of the autocorrelation time of collisions increases the discounted reward
[…]
First, remember that, we can express the expected discounted reward, $\langle G\rangle$, as a sum over the expected discounted reward conditioned on the reward observed at zero-th step, $R_{0}$,
@@
[…]
@@
$\langle G|R_{0}=r\rangle$ $=\frac{1}{\log(\gamma)}\left[\frac{\langle\rho\rangle\left(r_{-}\left(\langle\rho\rangle-\tau_{-}\log(\gamma)\right)-\langle\rho\rangle+1\right)}{\tau_{-}\log(\gamma)-1}-\frac{(\langle\rho\rangle-1)\left(\langle\rho\rangle r_{-}-\tau_{+}\log(\gamma)-\langle\rho\rangle+1\right)}{\tau_{+}\log(\gamma)-1}\right]$

### v2

@@ Results
A large number of environments tend to follow similar trends, indicating that the DNNs in particles do not distinguish between these environments in terms of repulsiveness and that DNNs initially do not interpret the full complexity of the environment. At characteristic times that partially overlap with the boundaries of the temporal regimes defined on the macroscopic scale, we observe branching points after which environments become distinguishable by the DNNs to increasing degrees. Therefore, over time, DNNs of particles learn to distinguish between an increasing number of environments at characteristic time points. In between these time points, the form of interactions tends to be stable over increasingly long time periods.
@@
Figure 7: To understand how interactions involve over time, we computed the time evolution of the repulsiveness, averaged over all particles. Each line shows the repulsiveness in a different environment. The arrow and the dashed circles indicate the characteristic branching points of the interactions. (a) Repulsiveness colored by the distance to the particle was added to the base environments. (b) Repulsiveness for the case that a particle is added at a distance $\Delta x=1$ to the base environments. Environments are coloured by the distance to the second nearest neighbour. (c) Repulsiveness for the case that a particle is added at a distance $\Delta x=1$ to the base environments. Environments are coloured by the distance to the second nearest neighbour.
@@
We next address the origin of the abrupt change in the behaviour of the particle as a function of the particle density and the associated decrease in rewards? Our analyses have shown both macroscopic order in space and time, which indicates collective behaviour of particles, and highly complex, density-dependent interactions that depend on evolving particle “species”. To test the hypothesis that the abrupt change in the behaviour of particles is a consequence of self-organisation, we now develop a minimal, phenomenological model that comprises only a minimal set of interactions without explicit density dependence.
To this end, we use the framework of active matter theory and Langevin equations. We consider the position $x_{i}(t)$, and velocity $v_{i}(t)$ of a particle $i$ at a given instance of time, $t$. The time evolution of the position and velocity of this particle is then described by a set of Langevin equations of the form
@@
[…]
@@
where $\tau_{-}$ and $\tau_{+}$ are the autocorrelation times of collisions and non-collisions, respectively. The minimisation of the autocorrelation time is empirically supported by our simulations (Fig. S1). The autocorrelation time is proportional to the time a pair of particles spends in each other’s vicinity, such that particles minimise a function that is proportional to 1/|vu|1/|v-u|. This then leads to a force that is given by the derivative of this term with respect to $v$, $\operatorname{sign}(v-u)/|v-u|^{2}$. Because the correlation scales with the density, the second contribution to the interaction terms takes the form $\rho\operatorname{sign}(v-u)/|v-u|^{2}$. Taken together, in our phenomenological description, this gives an effective force term for particle $i$ of the form
[…]
Figure 8: a) Representative trajectories from Langevin simulations of the equation (7) and (8), showing the positions of particles at the specified particle density in the last 3000 time points of the simulations, see Appendix B for details. The colours show the sign of the average velocity of particles in the last 100 time points (red: positive, blue: negative). For densities larger than 0.2, we have randomly sampled 256 particles to make individual lines distinguishable b) Phase portrait of Eqs. (9) and (10) for density 0.1. c) Phase portrait of Eqs. (9) and (10) for a density 0.5. The colours show the sign of the velocity of the test particle in the Hartree approximation. d) Velocity autocorrelation time obtained from simulations of Eqs. (7) and (8). e) Velocity autocorrelation time obtained from simulation of the full model in Fig. 1(b), see Appendix F for details. The dashed line shows the density of 0.3
We now define the active friction term $\gamma(v)$. Because our system is out of equilibrium, the fluctuation-dissipation theorem does not hold, such that friction is not coupled to the noise amplitude. Instead, this term comprises two parts: one that describes directed motion and another that describes friction. A phenomenological term that describes both is the Rayleigh-Helmholtz friction [50], which has been used to describe a variety of active matter systems. It is of the form
@@
where we have dropped the index $i$. We have non-dimensionalised the Langevin equations by rescaling time and space by $t^{\prime}=\alpha t$ and $x^{\prime}=x\sqrt{\alpha\beta}$, respectively, and removed the apostrophes to simplify the notation (see Appendix C). The dimensionless parameters $s=\zeta\beta^{2}/\alpha^{2}$ and $r=\delta\sqrt{\beta^{3}/\alpha}$ quantify the strength of interactions due to the minimisation of correlations and repulsion relative to the effect of self-propulsion, respectively. $B=D\beta/\alpha^{2}$ is the dimensionless noise strength.
In this model, for high densities, fluctuations in the particle density are weak, such that we expect the term that minimises correlations to dominate. Intuitively, this leads to a maximisation of the difference between the particle’s velocity $v$ and the average velocity $u$, which could globally lead to a split as observed in our simulations. For low densities, the term that minimises correlations is negligible, and we expect particle repulsion to be dominant.
To test this numerically, we performed stochastic simulations of Eqs. (7) and (8) using a finite difference method (see Appendix C). We indeed found that particles show global alignment at low densities and a split into groups moving in opposite directions for high densities (Fig. 8). Again, the variance of particles moving in the positive direction across simulations is smaller than expected by chance (t-test, $p=6.16e^{-7}$), suggesting a collective tuning of the sizes of both groups. These observations suggest that the abrupt change in the behaviour of particles, as seen in Fig. 3(a), is a phase transition resulting from collective behaviour resulting from the effective interactions between particles.
To investigate this more quantitatively, we analytically derived the coevolution of the velocity of a particle, $v$, and the average velocity of all particles $\bar{u}$. To this end, following Ref. [51], we used a moment expansion up to second order in the average velocity to obtain a set of nonlinear differential equations (see Appendix E),
$\dot{v}=$ $(1-v^{2})v+s\frac{\rho\operatorname{sign}{(v-\bar{u})}}{(v-\bar{u})^{2}}+r(\bar{u}-v)\,$, (9)
@@
Here, $T=\langle(v-\bar{u})^{2}\rangle$ is the variance of the velocities, which we estimated from our simulations of the Langevin equations. The co-evolution of the velocity of individual particles and the average velocity is summarised in phase portraits (Fig. 8). The phase portraits show attractors that correspond to asymptotic dynamical regimes of Eqs. (7) and (8). For low densities, we find two attractors with $v=\bar{u}$ that correspond to states where the velocity of particles aligns with the average velocity in the system. For high densities, these attractors are located horizontally on a line given by $\bar{u}=0$. Therefore, the phenomenological model predicts the evolution of individual velocities to take specific positive or negative values under the constraint of a vanishing global velocity. Because the line $\bar{u}=0$ is stable, fluctuations around it are suppressed by feedback mechanisms. This corresponds to the observed state with precisely tuned equal proportions of particles moving in either direction.
Taken together, our analysis suggests that the abrupt change in the behaviour as a function of the particle density is an emergent phenomenon resulting from the effective interactions between particles. It is therefore a phase transition as in statistical physics systems. At all densities, interactions between particles have a macroscopic effect. In the high-density regime, particles optimise their reward locally, without being able to predict the suboptimal consequences of their behaviour on the macroscopic scale. This resembles a learning analogue to frustration phenomena in disordered systems or the fallacy of the commons in the social sciences. This analysis raises the question of whether this phase transition resembles known phase transitions in (non-equilibrium) statistical physics. To investigate this, we calculated autocorrelation functions of the velocities of individual particles, both in the Langevin simulations and in the simulations of the full model comprising intelligent agents, for different particle densities. In both cases, the autocorrelation time diverges at the density at which the phase transition occurs (Fig. 8) . Diverging correlations are characteristic of critical points, suggesting critical behaviour. A further characterisation of the phase transition requires the computation of critical exponents, which is currently not feasible for intelligent agents. It is, however, of note that the phase transition occurs at a density at which the range covered by the repulsive interactions of each particle multiplied by the number of particles equals the lattice size; that is, the effective volume fraction is equal to 1. This could indicate a relation to jamming transitions.
@@ Discussion
Taken together, we studied a paradigmatic model of robotic matter. In this model, particles in a stochastic lattice gas are endowed with deep neural networks that take the spatial environment of a particle as input and predict transition rates. Our extensive simulations reveal a rich phenomenology, including complex emergent phenomena leading to a phase transition, the evolution of multiple particle species, temporal learning regimes, and a phenomenon akin to frustration, where particles adopt suboptimal strategies. The phase transition we observe shows signs of criticality. Although it is not possible to determine critical exponents, the behaviour close to the phase-transition point shows similarities to a jamming transition.
@@
Stochastic many-particle systems, such as the one studied in this work, differ qualitatively from systems typically studied in active matter theory: Because each particle has many more intrinsic degrees of freedom than the number of particles that one can typically simulate on computers, they effectively lack microscopic symmetries. While in statistical physics one is typically interested in the breaking of microscopic symmetries, here the question arises whether symmetries emerge as a consequence of learning, and whether these symmetries are then broken again. Because deep neural networks of sufficient size can, in principle, approximate arbitrarily complex function interactions between particles, they do not necessarily follow simple patterns. They may be qualitatively different between particles and environments. This raises the question of how far physical theories are adequate in describing the time evolution of such systems. In our work, particles indeed developed complex interactions that cannot be described by simple mathematical functions. However, a simple phenomenological theory focusing on the rudimentary aspects of these interactions can predict a key aspect of the system, namely the existence of a density-dependent phase transition.
Reinforcement learning has been successfully used to control the behaviour of non-equilibrium systems, by controlling the motion of a single particle subject to hydrodynamic forces [20, 52], by exerting global control on many-particle systems [53, 54], or by enforcing alignment by rewarding high particle densities [29]. In our work, the use of neural networks allows for the learning of almost arbitrarily complex relations between the environment of a particle and its behaviour. These possibilities are explored and used by the particles, as evidenced in the learning of increasingly complex interactions and the formation of particle species. This again leads to remarkably rich spatio-temporal behaviour.
@@
Although our model is a drastic simplification of real robotic systems, our results show similarities to much more complex robotic systems. For example, Ref. [37] studied a hide-and-seek game between intelligent agents, in which agents passed through different, metastable learning regimes, and interactions between agents led to a decrease in the collected reward. In addition to interacting robots, virtual artificial intelligence systems, such as widely used large language systems, are now beginning to be combined to achieve better performance [55]. This raises the question of how to design these systems to achieve an expected outcome. In some cases, learning objectives straightforwardly translate into collective behaviour [29]. Our work implies that learning can be drastically influenced in unpredictable ways by collective behaviour on the macroscopic scale.
@@ Acknowledgements
A.A. and O.B. contributed equally. A.A., O.B., and S.R. conceptualised the work. A.A. and S.R. supervised the work; O.B. wrote the simulation code and analysed data. O.B. performed analytical calculations. S.R. acquired funding. O.B. and S.R. wrote the initial draft of the manuscript. All authors edited and approved the final version of the manuscript.
DATA AVAILABILITY
The data are available from the authors upon reasonable request.
@@ A Implementation of the stochastic simulations
[…]
Where $n_{j}$ is the number of transitions the particle $j$ made during the episode, and $\delta_{i}:=R_{i}+\gamma V_{i+1}-V_{i}$, and $p_{i},R_{i},V_{i}$ are the probability with which the particle made the transition, the reward it got and the value predicted by the critic at step $i$, respectively, where $i$ run from $1$ until $(n_{j}-1)$. To train the critic, we used the loss $\sum_{i=1}^{n_{j}-1}\delta_{i}^{2}/(n_{j}-1)$.
Figure S1: The rescaled average autocorrelation time of collisions, averaged over $10$ replicates. The autocorrelation time for each density is rescaled by their value at episode 11. The shaded area around the curves denotes the standard error. The averages are calculated at logarithmically spaced time points.
@@ B Numerical solution of the Langevin equations
We used the Euler–Maruyama scheme to integrate the non-dimensionalised Langevin equation. We tested $N=200$ and $N=1024$ particles with time increment $\delta t=0.001/\tau_{t}$, where $\tau_{t}=1/\alpha$ is the characteristic time-scale, and for a total time $T=1\times 10^{5}/\tau_{t}$. The system size was determined based on $L=N/\langle\rho\rangle$ where $\langle\rho\rangle=\rho_{0}/\tau_{l}$ is the average non-dimensional particle density and $\tau_{l}=1/\sqrt{\alpha\beta}$ is the characteristic length-scale. We used periodic boundary conditions. We set $\alpha=16,\beta=1,\delta=50,\zeta=75,D=2.5$, and we set the radius of interactions, $R$, to $1/\tau_{l}$ to calculate the local density and velocity field $\rho(x),u(x)$. For $N=200$, we run $15$ to $24$ replicates per particle density for $N=200$, while for $N=1024$, we run 2 replicates per particle density due to the limited availability of GPU hardware. To prevent divergences in the numerical integration of the singular term $\frac{\operatorname{sign}{(v^{\prime}-u^{\prime})}}{(v^{\prime}-u^{\prime})^{2}}$, we regularised it with the $\tanh$ function.
@@ C Non-dimensionalisation of the Langevin equation
To non-dimensionalise the Langevin equations, we rescaled time and positions as
@@ D Minimisation of the autocorrelation time of collisions increases the discounted reward
In this appendix, we show that the expected discounted reward, $\langle G\rangle$, increases as the autocorrelation time of the collisions $\tau_{-}=\int_{0}^{\infty}\langle\delta(R_{0}-r_{-})\cdot\delta(R_{t}-r_{-})\rangle\differential t d$ecreases.
First, remember that we can express the expected discounted reward, $\langle G\rangle$, as a sum over the expected discounted reward conditioned on the reward observed at the zero-th step, $R_{0}$,
@@
$\langle G|R_{0}=r\rangle$ $=\int_{0}^{\infty}dt\gamma^{t}\langle R_{t}|R_{0}=r\rangle$
$=\int_{0}^{\infty}dt\gamma^{t}\left(\langle R_{t}\rangle+e^{-t/\tau}(r-\langle R_{0}\rangle)\right),$ (S.5)
@@
$\langle G\rangle$ $=\frac{1}{\log(\gamma)}\left[\frac{\langle\rho\rangle\left(r_{-}\left(\langle\rho\rangle-\tau_{-}\log(\gamma)\right)-\langle\rho\rangle+1\right)}{\tau_{-}\log(\gamma)-1}-\frac{(\langle\rho\rangle-1)\left(\langle\rho\rangle r_{-}-\tau_{+}\log(\gamma)-\langle\rho\rangle+1\right)}{\tau_{+}\log(\gamma)-1}\right]$
@@
E Coevolution of single-particle velocity $v$ and the average velocity $\bar{u}$
We follow the approach taken by references [51, 56] and use the moment-expansion method to find a closed set of equations for the moments of the probability density $P(x,v,t)$ of finding a particle at position $x$ with velocity $v$ at time $t$. We then truncate this hierarchy of equations at order two to arrive at an approximation of the true probability density. First, to be able to obtain a closed set of equations, we start by replacing the term $r\partial_{x}\rho$ in equation (8) with the term r(uv)r(u-v). This replacement allows us to express $\dot{v}$ in terms only of $v$ and $u$. Next, we define our moment-expansion hierarchy. Recall that, n-th moment, $M_{n}(x,t)$,
$M_{n}(x,t):=\left\langle v^{n}\right\rangle=\frac{1}{\rho}\int v^{n}P(x,v,t)dv,\quad n>0$ (S.7)
The normalisation $\rho$ is the zeroth moment, which is equivalent to the marginal density
$M_{0}(x,t)=\rho(x,t)=\int P(x,v,t)dv$ (S.8)
Multiplying the nth moment by the density and taking the derivative with respect to time, we arrive at the dynamics of the nth moments of the velocity
$\frac{\partial}{\partial t}\left(M_{0}M_{n}\right)=\int v^{n}\frac{\partial P}{\partial t}dv.$ (S.9)
Next, we use an effective single-particle description and use the Fokker-Planck equation to obtain
$\frac{\partial P}{\partial t}=$ $-v\frac{\partial}{\partial x}P+B\frac{\partial^{2}}{\partial v^{2}}P$ (S.10)
$-\frac{\partial}{\partial v}\left\{\left(1-v^{2}\right)v+s\frac{\rho\operatorname{sign}{(v-u)}}{(v-u)^{2}}+r\left(u-v\right)\right\}P$
Then Inserting equation (S.10) into equation (S.9), and using $\lim_{v\to\pm\infty}P(x,v,t)=0$, the terms containing partial derivatives with respect to $v$ can be partially integrated, resulting in
$\frac{\partial}{\partial t}\left(M_{0}M_{n}\right)=$ $-\frac{\partial}{\partial x}\rho\left\langle v^{n+1}\right\rangle+n(n-1)B\rho\left\langle v^{n-2}\right\rangle$ (S.11)
$+n\rho\left(\left\langle v^{n}\right\rangle-\left\langle v^{n+2}\right\rangle\right)$
$+n\rho\left(s\rho\int v^{n-1}\frac{v-u}{|v-u|^{3}}Pdv\right)$
$+n\rho\left(r\left\langle v^{n-1}\right\rangle-\left\langle v^{n}\right\rangle\right).$
We now express $v(x,t)=\bar{u}(x,t)+\delta v(x,t)$, as the average velocity plus deviations from it. We assume that $\langle\delta v^{l}\rangle=0$ for odd exponents $l=1,3,5,$... . Thus we obtain
$\langle v\rangle$ $=\bar{u}$ (S.12)
$\langle v^{2}\rangle$ $=\bar{u}^{2}+T$ (S.13)
$\langle\delta v^{2}\rangle$ $=T.$ (S.14)
The dynamics of the first moment are given by
$\dot{\bar{u}}=\bar{u}-\bar{u}(\bar{u}^{2}+3T)$ (S.15)
In the limit where the number of particles goes to infinity, one can approximate $u(x,t)$ with $\bar{u}(x,t)$. Furthermore, thanks to the translational invariance of the system, we can treat the fields $v(x,t),\bar{u}(x,t)$ as scalar variables $v(t),u(t)$. Now, embedding a tracer particle whose dynamics are given by equation (8) into a system whose dynamics are determined by equation (S.15), we obtain the following set of two equations.
$\dot{\bar{u}}$ $=\bar{u}-\bar{u}(\bar{u}^{2}+3T)$ (S.16)
$\dot{v}$ $=v-v^{3}+\zeta\frac{\beta^{2}}{\alpha^{2}}\frac{\rho\operatorname{sign}{(v-\bar{u})}}{(v-\bar{u})^{2}}+\delta\sqrt{\frac{1}{\alpha}}(\bar{u}-v)$ (S.17)
which qualitatively capture the phenomenology of the system with varying density.
Ideally, one can truncate the hierarchy at order three and above to obtain the dynamics of $T$. However, it was not possible to express the integral $\int v^{n-1}\frac{\operatorname{sign}{(v-\bar{u})}}{(v-\bar{u})^{2}}Pdv$ in terms of the moments of $v$ for $n>1$ to close to hierarchy. As a result, we measured the value of $T$ at different densities from the agent-based simulations and used those values in Fig. 8.
F Calculation of the velocity autocorrelation function
To calculate the velocity autocorrelation of individual particles in the full model comprising intelligent agents (Fig. 8(e)), we used the batch means method [57]. To calculate the velocity autocorrelation of individual particles in the many-body Langevin equation, we calculated the connected correlation function
$=\langle v_{i}(t+\Delta t)-v_{i}(t)\rangle-\langle v_{i}(t)\rangle\langle v_{i}(t+\Delta t)\rangle$ (S.18)
for each particle $i$. We then calculated the autocorrelation time using its definition
$\tau_{i}=\int_{0}^{T}d(\Delta t)C_{i}(\Delta t)/C_{i}(0).$ (S.19)
where $T$ was set to $500$.
G Gradient of the neural networks as a proxy for interactions
To understand how the dynamics of the DNN parameters relate to the formation of spatio-temporal structures, we analyse the forces the DNNs encode. Intuitively, interactions between particles are encoded in the temporal change in their velocity, averaged over all environments. To develop this idea, we first represent a given environment by a vector $\vec{e}$, which contains the number of particles at each position in the environment. To quantify the change in the velocity due to changes in the particle’s environment, we note that the change in the velocity, namely the force, can be written as
$F$ $=\langle\dot{v}\rangle$
[…]
Using the conservation of probability, we can replace $\partial_{t}P(\vec{e},t)$ with the divergence of the probability flux, $j$, giving us
$F$ $=-\int_{[0,\infty)^{l}}d\vec{e}(2p(\vec{e})-1)\nabla\cdot\vec{j}.$
Now, using integration by parts, we obtain
[…]
The first term in the integral can be turned into a surface integral via the divergence theorem, giving us
[…]
$=0,$ (S.22)
since, on the boundaries of the domain $0,\infty)^{l}$, the term $((2p(\vec{e})-1)\vec{j})\cdot\hat{n}$ vanishes either because $\vec{j}\cdot\hat{n}$ vanishes (in the case when all except one of $e_{i}$ is zero) or $(2p(\vec{e})-1$ goes to zero as $|\vec{e}|\to\infty$. As a result, finally we obtain
[…]
This motivates the use $\nabla p(\vec{e})$ as a proxy for the interactions of the particles.